# Cohomological Calibration and Curvature Constraints on Product Manifolds: A Topological Lower Bound

### v4

@@ 5. Main Result and Proof
Since $\nabla^{C}$ is metric, any non-trivial off-diagonal holonomy implies that $\mathfrak{hol}_{p}(\nabla^{C})$ acts irreducibly on $T_{p}M$. Theorem 5.2 quantifies the “forced irreducibility” principle in terms of the mixed tensor rank, modulo the parallel subspace $\mathcal{P}_{2}\otimes H^{1}(M_{2})$. Specifically, the dimension of the off-diagonal holonomy subspace is bounded below by
$\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\geq r^{\sharp}=r-\dim(\mathcal{P}_{2}\otimes H^{1}(M_{2})),$
where $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$. Irreducibility is forced if and only if $r^{\sharp}>0$, i.e. when the mixed tensor rank exceeds the dimension of the parallel subspace.
Remark 5.6 (Analyticity as a proof device).
The use of real-analytic metrics in the proof of Theorem 5.2 is purely instrumental: analytic approximation (Greene–Jacobowitz) and lower-semicontinuity of matrix rank extend the bound to every $C^{\infty}$ product metric without additional assumptions.
6. Examples and Sharpness
The case $M=S^{2}\times\Sigma_{g}$
Let $\Sigma_{g}$ be a closed oriented surface of genus $g\geq 1$. By Künneth,
$H^{3}(S^{2}\times\Sigma_{g};\mathbb{R})\cong H^{2}(S^{2})\otimes H^{1}(\Sigma_{g}),$ (6.1)
since $H^{1}(S^{2})=0$ and $H^{2}(\Sigma_{g})\cong\mathbb{R}$. The mixed component is therefore isomorphic to $\mathbb{R}^{2g}$. Choosing a harmonic basis $\{\alpha_{j}\}_{j=1}^{2g}$ for $H^{1}(\Sigma_{g})$ and the volume form $\mathrm{vol}_{S^{2}}$ on $S^{2}$, any mixed class has a harmonic representative of the form
$\omega_{h}=\sum_{j=1}^{2g}c_{j}\,\mathrm{vol}_{S^{2}}\wedge\alpha_{j}.$ (6.2)
Thus $\mathrm{rank}_{\mathbb{R}}\big([\omega]_{\mathrm{mixed}}\big)=2g$. For each $j$, select $X,Y\in V_{1}$ so that $\mathrm{vol}_{S^{2}}(X,Y)\neq 0$ and $Z\in V_{2}$ so that $\alpha_{j}(Z)\neq 0$; the corresponding curvature endomorphisms produce $2g$ independent off-diagonal directions. Hence
$\dim\big(\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\big)\ \geq\ 2g-\dim(\mathcal{P}_{2}\otimes H^{1}(\Sigma_{g})),$ (6.3)
where $\mathcal{P}_{2}\subset H^{2}(S^{2};\mathbb{R})$ is the subspace of harmonic parallel $2$-forms on $S^{2}$ (here $\mathcal{P}_{2}=\{0\}$, so the bound is $2g$).
The case $M=S^{2}\times T^{2}$
Here $H^{3}(M)\cong H^{2}(S^{2})\otimes H^{1}(T^{2})\cong\mathbb{R}^{2}$. Writing $\omega_{h}=c_{1}\,\mathrm{vol}_{S^{2}}\wedge dx+c_{2}\,\mathrm{vol}_{S^{2}}\wedge dy$ with $dx,dy$ the harmonic $1$-forms on the torus, one obtains two independent mixed components, hence
@@ 6. Examples and Sharpness
$\dim\big(\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\big)\ \geq\ 2-\dim(\mathcal{P}_{2}\otimes H^{1}(T^{2})),$ (6.4)
where $\mathcal{P}_{2}=\{0\}$ on $S^{2}$, so the bound is $2$.
@@ 7. Concluding Remarks and Perspectives
We have shown that for every compact and orientable product manifold, endowed with any product metric $g$ and for every metric connection satisfying $\nabla^{\mathrm{LC}}T\neq 0$, with totally skew torsion calibrated by a mixed cohomology class, the off-diagonal holonomy dimension admits a quantified lower bound
$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}=r-\dim(\mathcal{P}_{2}\otimes H^{1}(M_{2})),$
where $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$ and $\mathcal{P}_{2}\subset H^{2}(M_{1};\mathbb{R})$ is the space of harmonic parallel $2$-forms on $M_{1}$. The resulting lower bound is metric-invariant modulo parallel-form preserving deformations, thus providing the first computable obstruction to reducible holonomy on product manifolds. This quantifies how the topological “mixing strength” of $\omega]$ forces geometric “entanglement” across the product splitting, modulo the parallel subspace. The examples above demonstrate sharpness in natural settings where the torsion equals the harmonic representative and the parallel subspace has minimal dimension. It is natural to conjecture that equality holds whenever the torsion is purely harmonic, whereas strict inequality may occur when the coexact or exact parts of $T$ generate additional independent off-diagonal curvature directions or when the parallel subspace has maximal dimension. A systematic classification of such cases is an interesting direction for future work. A further interesting direction for future work could be to explore the implications of these results in the context of deformations of complex structures, a field pioneered by the works of Kodaira [6] and Kodaira-Spencer [7].

### v9

@@ 5. Main Result and Proof
Since $\nabla^{C}$ is metric, any non‑trivial off‑diagonal holonomy implies that $\mathfrak{hol}_{p}(\nabla^{C})$ acts irreducibly on $T_{p}M$. Theorem 5.2 quantifies this “forced irreducibility” principle in terms of the mixed tensor rank, modulo the obstruction space
$\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p},$
where, in the notation of the theorem,
$\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),\qquad\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr).$
Specifically, the dimension of the off‑diagonal holonomy subspace satisfies
@@
$\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\geq\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K},$
and irreducibility is forced precisely when the right‑hand side is positive, i.e. when
$\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})>\dim\mathcal{K}.$
In other words, a non‑zero value of the corrected rank $r^{\sharp}=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K}$ guarantees that the holonomy representation cannot split according to the product decomposition $T_{p}M=V_{1}\oplus V_{2}$; the connection is geometrically “entangled” across the two factors.
Remark 5.6 (Analyticity as a tool of proof).
The use of real‑analytic metrics in the proof of Theorem 5.2 is purely instrumental. By the density theorem of Greene–Jacobowitz [4] and the upper semicontinuity of the dimension of the space of parallel forms, one can approximate any smooth product metric $g=g_{1}\oplus g_{2}$ in the $C^{\infty}$-topology by a sequence of real‑analytic product metrics $\{g^{(k)}\}$ satisfying $\dim\mathcal{P}_{\ell}(M_{i},g^{(k)})\leq\dim\mathcal{P}_{\ell}(M_{i},g)$ for $\ell=1,2$.
For each analytic metric $g^{(k)}$ the curvature tensor $R_{\mathrm{off},k}(s)$ depends analytically on the deformation parameter $s$ and on the point $p\in M$; consequently the rank of the off‑diagonal curvature is lower‑semicontinuous and can be controlled by the behaviour of analytic minors (Identity Principle).
Passing to the limit $k\to\infty$ via the lower‑semicontinuity of matrix rank with respect to uniform convergence [9] yields the desired bound for the original $C^{\infty}$ metric $g$. Thus the analytic approximation does not impose any restriction on the result—it merely provides a convenient technical tool for extending the estimate from a generic dense subset to the whole manifold.
6. Illustrative Examples
The case $\Sigma_{g}\times T^{2}$ (Example of a non-trivial lower bound)
Fix $g\geq 2$ and endow $\Sigma_{g}$ with any hyperbolic metric and $T^{2}$ with the flat metric. Harmonic spaces:
$H^{1}(T^{2})=\operatorname{span}\{dx,dy\},\qquad H^{2}(\Sigma_{g})=\operatorname{span}\{\operatorname{vol}_{\Sigma}\},$
and choose a non-zero harmonic $1$-form $\alpha\in H^{1}(\Sigma_{g})$ (such an $\alpha$ exists because $b_{1}(\Sigma_{g})=2g>0$).
For the hyperbolic metric on $\Sigma_{g}$ we have
$\mathcal{P}_{1}(\Sigma_{g})=0,\qquad\mathcal{P}_{2}(\Sigma_{g})=\operatorname{span}\{\operatorname{vol}_{\Sigma}\}\cong\mathbb{R},$
since the volume form is harmonic and parallel.
For the flat metric on $T^{2}$ all harmonic forms are parallel, hence
$\mathcal{P}_{1}(T^{2})=H^{1}(T^{2})\cong\mathbb{R}^{2},\qquad\mathcal{P}_{2}(T^{2})=H^{2}(T^{2})\cong\mathbb{R}.$
Consider the mixed cohomology class
$\omega]=\underbrace{\operatorname{vol}_{\Sigma}\otimes dx}_{\in H^{2}(\Sigma_{g})\otimes H^{1}(T^{2})}\;+\;\underbrace{\alpha\otimes\operatorname{vol}_{T^{2}}}_{\in H^{1}(\Sigma_{g})\otimes H^{2}(T^{2})}\;\in\;H^{3}(\Sigma_{g}\times T^{2}).$
Both Künneth components are already simple tensors, therefore their minimal tensor ranks are
$r_{2,1}=1,\qquad r_{1,2}=1,$
and the mixed rank equals
$\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)=r_{2,1}+r_{1,2}=2.$
The obstruction kernels are computed as follows. For the $(2,1)$-part $\pi_{2,1}([\omega])=\operatorname{vol}_{\Sigma}\otimes dx$,
$\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr)=\operatorname{span}\{\operatorname{vol}_{\Sigma}\otimes dx\},$
because $\operatorname{vol}_{\Sigma}\in\mathcal{P}_{2}(\Sigma_{g})$ and $dx\in\mathcal{P}_{1}(T^{2})$. Hence $\dim\ker\Psi_{p}=1$.
For the $(1,2)$-part $\pi_{1,2}([\omega])=\alpha\otimes\operatorname{vol}_{T^{2}}$,
$\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr)=\{0\},$
because $\mathcal{P}_{1}(M_{1})=\{0\}$ and $\mathcal{P}_{2}(M_{2})=\operatorname{span}\{\operatorname{vol}_{T^{2}}\}$, hence $\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})=\{0\}$. Hence $\dim\ker\widetilde{\Psi}_{p}=0$. Thus
$\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p},\qquad\dim\mathcal{K}=1+0=1,$
and the corrected rank is
$r^{\sharp}=\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K}=2-1=1.$
Theorem 5.2 therefore yields the non‑trivial lower bound
$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\;\geq\;1.$
In particular, any cohomologically calibrated metric connection with non‑parallel torsion representing $\omega]$ necessarily has irreducible holonomy across the product splitting $\Sigma_{g}\times T^{2}$.
The case $T^{2}\times T^{2}$ (Example of a trivial lower bound)
Let $M=T^{2}\times T^{2}$ endowed with the flat product metric $g=g_{1}\oplus g_{2}$. Fix harmonic bases
$H^{1}(T^{2})=\operatorname{span}\{dx,dy\},\quad H^{2}(T^{2})=\operatorname{span}\{vol\}.$
For the flat metric we have
$\mathcal{P}_{1}(T^{2})=H^{1}(T^{2})\cong\mathbb{R}^{2},\qquad\mathcal{P}_{2}(T^{2})=H^{2}(T^{2})\cong\mathbb{R}.$
The mixed Künneth components are
$H^{2}(T^{2})\otimes H^{1}(T^{2})\cong\mathbb{R}^{2},\quad H^{1}(T^{2})\otimes H^{2}(T^{2})\cong\mathbb{R}^{2}.$
Consider the single cohomology class
$\omega]=\underbrace{vol\otimes dx}_{(2,1)}+\underbrace{dy\otimes vol}_{(1,2)}.$
Its minimal tensor-rank decomposition is
$\pi_{2,1}([\omega])=vol\otimes dx\quad(\text{rank }1),\qquad\pi_{1,2}([\omega])=dy\otimes vol\quad(\text{rank }1).$
Hence
$\operatorname{rank}_{\mathbb{R}}([\omega]_{\text{mixed}})=1+1=2.$
Since $vol\in\mathcal{P}_{2}$ and $dx,dy\in\mathcal{P}_{1}$, both simple tensors lie in the obstruction spaces
$\ker\Psi_{p}=\operatorname{span}\{vol\otimes dx\},\qquad\ker\tilde{\Psi}_{p}=\operatorname{span}\{dy\otimes vol\},$
so $\dim\mathcal{K}=2$. Theorem 5.2 gives
$\dim\mathfrak{hol}^{\text{off}}_{p}(\nabla^{C})\geq 2-2=0.$
7. Significance and Applications to Physics and STEM Disciplines
Connections with torsion and product structures appear naturally in a wide range of mathematical and physical models. In string theory, supergravity, and related frameworks, metric connections with totally skew-symmetric torsion play a central role in flux compactifications, generalized geometries, and effective field theories on product manifolds such as $M_{4}\times X$. In these settings, the holonomy of the canonical connection governs the coupling between the external (space-time) and internal (compactified) factors, and irreducibility prevents the decoupling of physical degrees of freedom.
Our lower bound provides an efficient mechanism for establishing such irreducibility directly from global invariants - particularly the mixed tensor rank and the corrected quantity $r^{\sharp}$ - without computing curvature tensors or higher covariant derivatives. This is especially relevant in high-dimensional backgrounds where explicit curvature calculations are notoriously difficult or mathematically intractable.
A similar phenomenon arises in gauge theory, condensed matter, and materials science, where product manifolds model layered, composite, or multi-component systems. The holonomy algebra encodes the geometric interactions between these components; detecting whether such interactions are essential (irreducible) or merely apparent becomes a fundamental structural question. The criterion developed in this paper identifies irreducibility based solely on the existence and non-parallelity of mixed differential forms, providing a geometric diagnostic that is robust under perturbations of the underlying metric and torsion.
Applications also extend to geometric analysis and computational modeling. In multi-scale PDE systems defined on product domain (such as transport-diffusion equations, elasticity models, or geometric flows) the holonomy of a natural geometric connection determines how different subsystems interact and whether the governing equations admit a decoupled formulation. Our lower bound offers a rigorous obstruction to such decoupling, enabling researchers to understand when full coupling is intrinsic to the geometry. This is particularly useful in numerical simulations, where direct computation of curvatures is expensive and unstable, while global invariants such as parallel-form spaces are straightforward to detect algorithmically.
Overall, the ability to certify irreducible holonomy through topological or geometric invariants provides a powerful tool across physics, geometry, and applied sciences. It eliminates the need for explicit curvature computations and transforms a classically analytic problem into a global geometric one, opening the door to applications in any setting where torsion, product structures, and holonomy play a foundational role.
@@ 8. Concluding Remarks and Perspectives
We have shown that for every compact and orientable product manifold, endowed with any product metric $g$ and for every metric connection satisfying $\nabla^{\mathrm{LC}}T\neq 0$, with 3-form torsion representing a mixed cohomology class via its harmonic projection $\omega_{h}$, the off-diagonal holonomy dimension admits a quantified lower bound
$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}=r-\dim\mathcal{K},$
where $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$ and $\mathcal{K}=\ker\Psi_{p}\oplus\ker\tilde{\Psi}_{p}$ is the sum of the kernels of the linear maps $\Psi_{p},\tilde{\Psi}_{p}$ that encode the obstruction to off-diagonal curvature. The resulting lower bound is metric-invariant modulo parallel-form preserving deformations, thus providing the first computable obstruction to reducible holonomy on product manifolds. This quantifies how the topological “mixing strength” of $\omega]$ forces geometric “entanglement” across the product splitting, modulo the parallel subspace contained within the mixed class. The examples above illustrate the computation of $r^{\sharp}$ in simple geometric settings. When the torsion is purely harmonic and the parallel-form spaces are minimal, the bound is often saturated in explicit constructions (e.g., on flat tori with parallel torsion). It is natural to conjecture that equality holds generically for purely harmonic torsion, while strict inequality may occur when the coexact or exact parts of $T$ generate additional off-diagonal curvature directions, or when the parallel-form spaces are larger. A systematic classification of the cases where the bound is sharp remains an interesting direction for future work.
A further interesting direction for future work could be to explore the implications of these results in the context of deformations of complex structures, a field pioneered by the works of Kodaira [6] and Kodaira-Spencer [7].